# Frequency-Guided Masking for Enhanced Vision Self-Supervised Learning

### v1

@@ B.3 Ablation Study - Different Filters
Additionally, our masking approach generates unique masks according to each image’s frequency responses, thereby accounting for each image’s distinctive features and semantics (see examples in the Tables 9, 10 and 11). This contrasts sharply with random masking. By targeting essential frequencies, selective masking ensures that the model adapts to recognize and prioritize these key signals, resulting in more robust and effective pre-training. Our findings indicate that this method significantly enhances the model’s generalization capabilities and overall accuracy, confirming the efficacy of selective masking in the frequency domain for developing advanced predictive models.
@@
Figure 4 illustrates the implementation of three supplementary random filtering techniques on the input frequencies and the effect of those filters compared with our filter. The filter that we directly applied in the frequency domain, [Kamarainen et al., 2006]. In addition to this, is utilized, drawing inspiration from the approach [Xie et al., 2022], where random square regions are masked in the frequency domain to mimic missing information. Furthermore, we introduce the strategy, which involves randomly masking circular areas at various distances from the center of the frequency spectrum. One of the significant advantages of our introduced filter (/) is that it uses actual image information for masking, while other filters just randomly or constantly mask (like MFM) a portion of the frequency area without considering the structure and information of the image.
@@
B.4 Efficiency Analysis
@@
To assess the effectiveness of our enhanced model, we provide memory usage on the GPU, and overall accuracy. The results of these evaluations are summarized in Table 8.
@@
Table 8 summarizes the performance metrics of various self-supervised learning methods, highlighting the impact of integrating dual inputs—original and filtered images—on model efficacy, particularly within the FOLK framework. Memory usage is a crucial consideration, as only GPUs with high capacity can run the model. This is noteworthy when compared to iBOT and AttMask which consume more memory222AttMask cannot run with a batch size of 128 on an A100 80GB GPU, so we run it with a batch size of 64., while our model requires less than 20GB of memory for the same batch size. The increased memory usage in models like FOLK and iBOT correlates with the advanced processing capabilities necessary for handling dual inputs. However, the evident gains in learning accuracy justify the resource allocation, making it a worthwhile trade-off for applications where high precision and robust feature recognition are critical, such as few-shot learning scenarios. FOLK achieves the highest few-shot learning accuracy at and tops classification accuracy at . This performance indicates that the model’s ability to effectively utilize both filtered and original images significantly enhances its learning capabilities, particularly in tasks requiring robust feature recognition.
Appendix C Prediction Visualization
@@
This section presents visualizations of a series of images from the ImageNet-1K dataset, processed by our newly proposed techniques, namely the and filters. Tables 9, 10, and 11 collectively demonstrate the impact of our proposed approach. Distinct from the MFM method, which employs static and conventional low/high-pass filters, and filters are dynamic and tailored to each image. This adaptability allows the filters to change in response to an image’s unique concept and structural characteristics, offering a more nuanced and effective processing method. In addition, Tables 9, 10, and 11 also show the reconstructed missing frequencies for each image, generated by the FOLK pre-trained model. The clear alignments between the model predicted and the ground-truth masked frequencies provide further evidence for the successful pre-training of FOLK and, therefore the improved model efficacy.
@@
![[Uncaptioned image]](extracted/5858280/images/1/org.png)
![[Uncaptioned image]](extracted/5858280/images/1/freq.png)
![[Uncaptioned image]](extracted/5858280/images/1/freq_50.png)
![[Uncaptioned image]](extracted/5858280/images/1/org_50.png)
![[Uncaptioned image]](extracted/5858280/images/1/_out_50.png)
![[Uncaptioned image]](extracted/5858280/images/1/freq_50_m.png)
![[Uncaptioned image]](extracted/5858280/images/1/org_50_m.png)
![[Uncaptioned image]](extracted/5858280/images/1/_out_50_m.png)
![[Uncaptioned image]](extracted/5858280/images/1/freq_10.png)
![[Uncaptioned image]](extracted/5858280/images/1/org_10.png)
![[Uncaptioned image]](extracted/5858280/images/1/_out_10.png)
![[Uncaptioned image]](extracted/5858280/images/1/freq_10_m.png)
![[Uncaptioned image]](extracted/5858280/images/1/org_10_m.png)
![[Uncaptioned image]](extracted/5858280/images/1/_out_10_m.png)
![[Uncaptioned image]](extracted/5858280/images/1/freq_5.png)
![[Uncaptioned image]](extracted/5858280/images/1/org_5.png)
![[Uncaptioned image]](extracted/5858280/images/1/_out_5.png)
![[Uncaptioned image]](extracted/5858280/images/1/freq_5_m.png)
![[Uncaptioned image]](extracted/5858280/images/1/org_5_m.png)
![[Uncaptioned image]](extracted/5858280/images/1/_out_5_m.png)
![[Uncaptioned image]](extracted/5858280/images/2/org.png)
![[Uncaptioned image]](extracted/5858280/images/2/freq.png)
![[Uncaptioned image]](extracted/5858280/images/2/freq_50.png)
![[Uncaptioned image]](extracted/5858280/images/2/org_50.png)
![[Uncaptioned image]](extracted/5858280/images/2/_out_50.png)
![[Uncaptioned image]](extracted/5858280/images/2/freq_50_m.png)
![[Uncaptioned image]](extracted/5858280/images/2/org_50_m.png)
![[Uncaptioned image]](extracted/5858280/images/2/_out_50_m.png)
![[Uncaptioned image]](extracted/5858280/images/2/freq_10.png)
![[Uncaptioned image]](extracted/5858280/images/2/org_10.png)
![[Uncaptioned image]](extracted/5858280/images/2/_out_10.png)
![[Uncaptioned image]](extracted/5858280/images/2/freq_10_m.png)
![[Uncaptioned image]](extracted/5858280/images/2/org_10_m.png)
![[Uncaptioned image]](extracted/5858280/images/2/_out_10_m.png)
![[Uncaptioned image]](extracted/5858280/images/2/freq_5.png)
![[Uncaptioned image]](extracted/5858280/images/2/org_5.png)
![[Uncaptioned image]](extracted/5858280/images/2/_out_5.png)
![[Uncaptioned image]](extracted/5858280/images/2/freq_5_m.png)
![[Uncaptioned image]](extracted/5858280/images/2/org_5_m.png)
![[Uncaptioned image]](extracted/5858280/images/2/_out_5_m.png)
![[Uncaptioned image]](extracted/5858280/images/3/org.png)
![[Uncaptioned image]](extracted/5858280/images/3/freq.png)
![[Uncaptioned image]](extracted/5858280/images/3/freq_50.png)
![[Uncaptioned image]](extracted/5858280/images/3/org_50.png)
![[Uncaptioned image]](extracted/5858280/images/3/_out_50.png)
![[Uncaptioned image]](extracted/5858280/images/3/freq_50_m.png)
![[Uncaptioned image]](extracted/5858280/images/3/org_50_m.png)
![[Uncaptioned image]](extracted/5858280/images/3/_out_50_m.png)
![[Uncaptioned image]](extracted/5858280/images/3/freq_10.png)
![[Uncaptioned image]](extracted/5858280/images/3/org_10.png)
![[Uncaptioned image]](extracted/5858280/images/3/_out_10.png)
![[Uncaptioned image]](extracted/5858280/images/3/freq_10_m.png)
![[Uncaptioned image]](extracted/5858280/images/3/org_10_m.png)
![[Uncaptioned image]](extracted/5858280/images/3/_out_10_m.png)
![[Uncaptioned image]](extracted/5858280/images/3/freq_5.png)
![[Uncaptioned image]](extracted/5858280/images/3/org_5.png)
![[Uncaptioned image]](extracted/5858280/images/3/_out_5.png)
![[Uncaptioned image]](extracted/5858280/images/3/freq_5_m.png)
![[Uncaptioned image]](extracted/5858280/images/3/org_5_m.png)
![[Uncaptioned image]](extracted/5858280/images/3/_out_5_m.png)
![[Uncaptioned image]](extracted/5858280/images/4/org.png)
![[Uncaptioned image]](extracted/5858280/images/4/freq.png)
![[Uncaptioned image]](extracted/5858280/images/4/freq_50.png)
![[Uncaptioned image]](extracted/5858280/images/4/org_50.png)
![[Uncaptioned image]](extracted/5858280/images/4/_out_50.png)
![[Uncaptioned image]](extracted/5858280/images/4/freq_50_m.png)
![[Uncaptioned image]](extracted/5858280/images/4/org_50_m.png)
![[Uncaptioned image]](extracted/5858280/images/4/_out_50_m.png)
![[Uncaptioned image]](extracted/5858280/images/4/freq_10.png)
![[Uncaptioned image]](extracted/5858280/images/4/org_10.png)
![[Uncaptioned image]](extracted/5858280/images/4/_out_10.png)
![[Uncaptioned image]](extracted/5858280/images/4/freq_10_m.png)
![[Uncaptioned image]](extracted/5858280/images/4/org_10_m.png)
![[Uncaptioned image]](extracted/5858280/images/4/_out_10_m.png)
![[Uncaptioned image]](extracted/5858280/images/4/freq_5.png)
![[Uncaptioned image]](extracted/5858280/images/4/org_5.png)
![[Uncaptioned image]](extracted/5858280/images/4/_out_5.png)
![[Uncaptioned image]](extracted/5858280/images/4/freq_5_m.png)
![[Uncaptioned image]](extracted/5858280/images/4/org_5_m.png)
![[Uncaptioned image]](extracted/5858280/images/4/_out_5_m.png)
![[Uncaptioned image]](extracted/5858280/images/5/org.png)
![[Uncaptioned image]](extracted/5858280/images/5/freq.png)
![[Uncaptioned image]](extracted/5858280/images/5/freq_50.png)
![[Uncaptioned image]](extracted/5858280/images/5/org_50.png)
![[Uncaptioned image]](extracted/5858280/images/5/_out_50.png)
![[Uncaptioned image]](extracted/5858280/images/5/freq_50_m.png)
![[Uncaptioned image]](extracted/5858280/images/5/org_50_m.png)
![[Uncaptioned image]](extracted/5858280/images/5/_out_50_m.png)
![[Uncaptioned image]](extracted/5858280/images/5/freq_10.png)
![[Uncaptioned image]](extracted/5858280/images/5/org_10.png)
![[Uncaptioned image]](extracted/5858280/images/5/_out_10.png)
![[Uncaptioned image]](extracted/5858280/images/5/freq_10_m.png)
![[Uncaptioned image]](extracted/5858280/images/5/org_10_m.png)
![[Uncaptioned image]](extracted/5858280/images/5/_out_10_m.png)
![[Uncaptioned image]](extracted/5858280/images/5/freq_5.png)
![[Uncaptioned image]](extracted/5858280/images/5/org_5.png)
![[Uncaptioned image]](extracted/5858280/images/5/_out_5.png)
![[Uncaptioned image]](extracted/5858280/images/5/freq_5_m.png)
![[Uncaptioned image]](extracted/5858280/images/5/org_5_m.png)
![[Uncaptioned image]](extracted/5858280/images/5/_out_5_m.png)
![[Uncaptioned image]](extracted/5858280/images/6/org.png)
![[Uncaptioned image]](extracted/5858280/images/6/freq.png)
![[Uncaptioned image]](extracted/5858280/images/6/freq_50.png)
![[Uncaptioned image]](extracted/5858280/images/6/org_50.png)
![[Uncaptioned image]](extracted/5858280/images/6/_out_50.png)
![[Uncaptioned image]](extracted/5858280/images/6/freq_50_m.png)
![[Uncaptioned image]](extracted/5858280/images/6/org_50_m.png)
![[Uncaptioned image]](extracted/5858280/images/6/_out_50_m.png)
![[Uncaptioned image]](extracted/5858280/images/6/freq_10.png)
![[Uncaptioned image]](extracted/5858280/images/6/org_10.png)
![[Uncaptioned image]](extracted/5858280/images/6/_out_10.png)
![[Uncaptioned image]](extracted/5858280/images/6/freq_10_m.png)
![[Uncaptioned image]](extracted/5858280/images/6/org_10_m.png)
![[Uncaptioned image]](extracted/5858280/images/6/_out_10_m.png)
![[Uncaptioned image]](extracted/5858280/images/6/freq_5.png)
![[Uncaptioned image]](extracted/5858280/images/6/org_5.png)
![[Uncaptioned image]](extracted/5858280/images/6/_out_5.png)
![[Uncaptioned image]](extracted/5858280/images/6/freq_5_m.png)
![[Uncaptioned image]](extracted/5858280/images/6/org_5_m.png)
![[Uncaptioned image]](extracted/5858280/images/6/_out_5_m.png)

### v2

@@ B.5.4 Different Filters
Additionally, our masking approach generates unique masks according to each image’s frequency responses, thereby accounting for each image’s distinctive features and semantics (see examples in the Tables 13, 14 and 15). This contrasts sharply with random masking. By targeting essential frequencies, selective masking ensures that the model adapts to recognize and prioritize these key signals, resulting in more robust and effective pre-training. Our findings indicate that this method significantly enhances the model’s generalization capabilities and overall accuracy, confirming the efficacy of selective masking in the frequency domain for developing advanced predictive models.
@@
Figure 4 illustrates the implementation of three supplementary random filtering techniques on the input frequencies and the effect of those filters compared with our filter. The filter that we directly applied in the frequency domain, (Kamarainen et al., 2006). In addition to this, is utilized, drawing inspiration from the approach (Xie et al., 2022), where random square regions are masked in the frequency domain to mimic missing information. Furthermore, we introduce the strategy, which involves randomly masking circular areas at various distances from the center of the frequency spectrum. One of the significant advantages of our introduced filter (/) is that it uses actual image information for masking, while other filters just randomly or constantly mask (like MFM) a portion of the frequency area without considering the structure and information of the image.
@@
B.5.5 Effect of Informed Filters on Pre-training
To further illustrate the impact of the proposed and filters on the pre-training process, we present in Fig. 5 a comparison of the loss progression of MFM when utilizing the original constant filters versus the / filters. Importantly, both experiments were conducted without the self-distillation design of FOLK, thereby isolating the analysis to the effect of the filters alone. As depicted in the figure, using the proposed and filters results in a higher MFM reconstruction loss (Eq. 3) throughout the pre-training process. This indicates that the model finds it more challenging to predict the masked frequencies produced by the / filters, suggesting that our filters are more effective in encouraging the model to learn richer representations by presenting more demanding reconstruction tasks.
B.6 Efficiency Analysis
@@
To assess the effectiveness of our enhanced model, we provide memory usage on the GPU, and overall accuracy. Table 12 summarizes the performance metrics of various self-supervised learning methods, highlighting the impact of integrating dual inputs—original and filtered images—on model efficacy, particularly within the FOLK framework. Memory usage is a crucial consideration, as only GPUs with high capacity can run the model. This is noteworthy when compared to iBOT and AttMask which consume more memory222AttMask cannot run with a batch size of 128 on an A100 80GB GPU, so we run it with a batch size of 64., while our model requires less than 20GB of memory for the same batch size. The increased memory usage in models like FOLK and iBOT correlates with the advanced processing capabilities necessary for handling dual inputs. However, the evident gains in learning accuracy justify the resource allocation, making it a worthwhile trade-off for applications where high precision and robust feature recognition are critical, such as few-shot learning scenarios. FOLK achieves the highest few-shot learning accuracy at and tops classification accuracy at . This performance indicates that the model’s ability to effectively utilize both filtered and original images significantly enhances its learning capabilities, particularly in tasks requiring robust feature recognition.
Appendix C Comparison with DINO and Distinctiveness of FOLK
While our proposed FOLK framework and DINO (Caron et al., 2021) both employ self-distillation techniques, it is crucial to clarify the key differences between the two methods and highlight the unique contributions of FOLK.
C.1 Use of Self-Distillation
Both FOLK and DINO utilize knowledge distillation in a self-supervised learning context involving a student-teacher architecture where the student model learns from the teacher model’s output. However, knowledge distillation is a common technique in self-supervised learning (SSL) and has been adopted in several works such as BYOL (Grill et al., 2020), iBOT (Zhou et al., 2022), and AttMask (Kakogeorgiou et al., 2022). Therefore, using self-distillation alone does not define a method’s identity or make it equivalent to DINO.
C.2 Addressing Different Problems
FOLK integrates self-distillation to address a specific limitation in Masked Frequency Modeling (MFM) (Xie et al., 2023). Namely, the model lacks looking at the natural view of images and unmasked images during pre-training. This limitation can hinder the model’s performance, especially in few-shot learning scenarios where labeled data is scarce. By incorporating self-distillation, FOLK allows the student model, which only sees frequency-masked images, to learn from the teacher model’s representations of unmasked images, thereby enhancing its ability to generalize to natural images.
C.3 Frequency-Based Masking and Filters
A key distinction of FOLK is the introduction of informed frequency-based filters ( and ) for masking in the frequency domain. These filters are designed to adaptively mask significant frequencies based on each image’s unique frequency components, creating a more challenging pretext task and enhancing the model’s ability to learn both macro and micro visual cues. This approach fundamentally differs from DINO, which operates in the spatial domain.
C.4 Computational Efficiency
FOLK is designed to be computationally efficient compared to DINO. Notably, FOLK requires fewer pre-training epochs to achieve competitive or superior performance (see Table 1). For instance, FOLK achieves strong results with 300 or 800 pre-training epochs, whereas DINO typically requires 800 to 1600 epochs to reach even lower performance levels (Caron et al., 2021).
Additionally, FOLK has lower GPU memory requirements. Using a batch size of 128, FOLK’s GPU usage is approximately 19.8 GB (see B.6), while DINO’s GPU usage is around 30.8 GB (based on reported usage in (Caron et al., 2021) with adjusted batch sizes). This efficiency is partly due to FOLK not requiring additional local views of images during training, unlike DINO.
C.5 Methodological Differences
While both methods use self-distillation, FOLK and DINO differ in their training methodologies:
Input Views: FOLK employs frequency-masked views generated using and filters, whereas DINO uses multiple spatial views, including global and local crops. FOLK does not need any local view.
Loss Functions: FOLK combines a reconstruction loss (MFM loss) for masked frequency prediction with a distillation loss, enabling the model to learn both from reconstructing missing frequencies and from the teacher’s representations. DINO relies solely on the distillation loss.
Exposure to Original Images: In FOLK, the teacher model is exposed to unmasked images, and the student learns to reconstruct the teacher’s representations from frequency-masked images. This design helps the student model adapt to natural images despite not seeing them directly during training.
Appendix D Prediction Visualization
@@
This section presents visualizations of a series of images from the ImageNet-1K dataset, processed by our newly proposed techniques, namely the and filters. Tables 13, 14, and 15 collectively demonstrate the impact of our proposed approach. Distinct from the MFM method, which employs static and conventional low/high-pass filters, and filters are dynamic and tailored to each image. This adaptability allows the filters to change in response to an image’s unique concept and structural characteristics, offering a more nuanced and effective processing method. In addition, Tables 13, 14, and 15 also show the reconstructed missing frequencies for each image, generated by the FOLK pre-trained model. The clear alignments between the model predicted and the ground-truth masked frequencies provide further evidence for the successful pre-training of FOLK and, therefore the improved model efficacy.
@@
![[Uncaptioned image]](extracted/6275800/images/1/org.png)
![[Uncaptioned image]](extracted/6275800/images/1/freq.png)
![[Uncaptioned image]](extracted/6275800/images/1/freq_50.png)
![[Uncaptioned image]](extracted/6275800/images/1/org_50.png)
![[Uncaptioned image]](extracted/6275800/images/1/_out_50.png)
![[Uncaptioned image]](extracted/6275800/images/1/freq_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/1/org_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/1/_out_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/1/freq_10.png)
![[Uncaptioned image]](extracted/6275800/images/1/org_10.png)
![[Uncaptioned image]](extracted/6275800/images/1/_out_10.png)
![[Uncaptioned image]](extracted/6275800/images/1/freq_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/1/org_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/1/_out_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/1/freq_5.png)
![[Uncaptioned image]](extracted/6275800/images/1/org_5.png)
![[Uncaptioned image]](extracted/6275800/images/1/_out_5.png)
![[Uncaptioned image]](extracted/6275800/images/1/freq_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/1/org_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/1/_out_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/2/org.png)
![[Uncaptioned image]](extracted/6275800/images/2/freq.png)
![[Uncaptioned image]](extracted/6275800/images/2/freq_50.png)
![[Uncaptioned image]](extracted/6275800/images/2/org_50.png)
![[Uncaptioned image]](extracted/6275800/images/2/_out_50.png)
![[Uncaptioned image]](extracted/6275800/images/2/freq_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/2/org_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/2/_out_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/2/freq_10.png)
![[Uncaptioned image]](extracted/6275800/images/2/org_10.png)
![[Uncaptioned image]](extracted/6275800/images/2/_out_10.png)
![[Uncaptioned image]](extracted/6275800/images/2/freq_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/2/org_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/2/_out_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/2/freq_5.png)
![[Uncaptioned image]](extracted/6275800/images/2/org_5.png)
![[Uncaptioned image]](extracted/6275800/images/2/_out_5.png)
![[Uncaptioned image]](extracted/6275800/images/2/freq_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/2/org_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/2/_out_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/3/org.png)
![[Uncaptioned image]](extracted/6275800/images/3/freq.png)
![[Uncaptioned image]](extracted/6275800/images/3/freq_50.png)
![[Uncaptioned image]](extracted/6275800/images/3/org_50.png)
![[Uncaptioned image]](extracted/6275800/images/3/_out_50.png)
![[Uncaptioned image]](extracted/6275800/images/3/freq_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/3/org_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/3/_out_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/3/freq_10.png)
![[Uncaptioned image]](extracted/6275800/images/3/org_10.png)
![[Uncaptioned image]](extracted/6275800/images/3/_out_10.png)
![[Uncaptioned image]](extracted/6275800/images/3/freq_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/3/org_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/3/_out_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/3/freq_5.png)
![[Uncaptioned image]](extracted/6275800/images/3/org_5.png)
![[Uncaptioned image]](extracted/6275800/images/3/_out_5.png)
![[Uncaptioned image]](extracted/6275800/images/3/freq_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/3/org_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/3/_out_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/4/org.png)
![[Uncaptioned image]](extracted/6275800/images/4/freq.png)
![[Uncaptioned image]](extracted/6275800/images/4/freq_50.png)
![[Uncaptioned image]](extracted/6275800/images/4/org_50.png)
![[Uncaptioned image]](extracted/6275800/images/4/_out_50.png)
![[Uncaptioned image]](extracted/6275800/images/4/freq_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/4/org_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/4/_out_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/4/freq_10.png)
![[Uncaptioned image]](extracted/6275800/images/4/org_10.png)
![[Uncaptioned image]](extracted/6275800/images/4/_out_10.png)
![[Uncaptioned image]](extracted/6275800/images/4/freq_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/4/org_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/4/_out_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/4/freq_5.png)
![[Uncaptioned image]](extracted/6275800/images/4/org_5.png)
![[Uncaptioned image]](extracted/6275800/images/4/_out_5.png)
![[Uncaptioned image]](extracted/6275800/images/4/freq_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/4/org_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/4/_out_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/5/org.png)
![[Uncaptioned image]](extracted/6275800/images/5/freq.png)
![[Uncaptioned image]](extracted/6275800/images/5/freq_50.png)
![[Uncaptioned image]](extracted/6275800/images/5/org_50.png)
![[Uncaptioned image]](extracted/6275800/images/5/_out_50.png)
![[Uncaptioned image]](extracted/6275800/images/5/freq_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/5/org_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/5/_out_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/5/freq_10.png)
![[Uncaptioned image]](extracted/6275800/images/5/org_10.png)
![[Uncaptioned image]](extracted/6275800/images/5/_out_10.png)
![[Uncaptioned image]](extracted/6275800/images/5/freq_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/5/org_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/5/_out_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/5/freq_5.png)
![[Uncaptioned image]](extracted/6275800/images/5/org_5.png)
![[Uncaptioned image]](extracted/6275800/images/5/_out_5.png)
![[Uncaptioned image]](extracted/6275800/images/5/freq_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/5/org_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/5/_out_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/6/org.png)
![[Uncaptioned image]](extracted/6275800/images/6/freq.png)
![[Uncaptioned image]](extracted/6275800/images/6/freq_50.png)
![[Uncaptioned image]](extracted/6275800/images/6/org_50.png)
![[Uncaptioned image]](extracted/6275800/images/6/_out_50.png)
![[Uncaptioned image]](extracted/6275800/images/6/freq_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/6/org_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/6/_out_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/6/freq_10.png)
![[Uncaptioned image]](extracted/6275800/images/6/org_10.png)
![[Uncaptioned image]](extracted/6275800/images/6/_out_10.png)
![[Uncaptioned image]](extracted/6275800/images/6/freq_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/6/org_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/6/_out_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/6/freq_5.png)
![[Uncaptioned image]](extracted/6275800/images/6/org_5.png)
![[Uncaptioned image]](extracted/6275800/images/6/_out_5.png)
![[Uncaptioned image]](extracted/6275800/images/6/freq_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/6/org_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/6/_out_5_m.png)